What force does the electric field of a charged infinite plane act with on each metre of a charged infinitely long filament placed in this field? The charge linear density on the filament is 3x10^- 8 C/cm and its surface density on the plane is 2x10^-8 C/cm2.



1
Expert's answer
2022-01-31T07:34:18-0500

"E=\frac{\sigma}{2\varepsilon_0},"

"E=\frac Fq,~q=\tau l,"

"\frac F{\tau l}=\frac{\sigma}{2\varepsilon_0},"

"F_1=\frac Fl=\frac{\sigma\tau}{2\varepsilon_0}=33.9~\frac Nm."